OpenStudy (anonymous):

Given k(x)=f(g(x)) and f(2)=-4 g(2)=2 f'(2)=3 g'(2)=5 Find k(2) and k'(2)

zepdrix (zepdrix):

\(\Large\bf \color{#CC0033 }{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf k(x)=f(g(x))\]By the chain rule:\[\Large\bf\sf k'(x)=f'(g(x))\cdot g'(x)\]

zepdrix (zepdrix):

\[\Large\bf\sf k'(2)=f'(g(2))\cdot g'(2)\]

zepdrix (zepdrix):

Then we just use the information they provided us with to plug in the pieces.

OpenStudy (anonymous):

That helps a little. Now I get: k(2) = -4 and k'(2) = - 8 How does that look?

zepdrix (zepdrix):

Your k(2) looks correct:\[\Large\bf\sf \color{royalblue}{g(2)=2}\]\[\Large\bf\sf f(2)=-4\] \[\Large\bf\sf k(2)\quad=\quad f\left[\color{royalblue}{g(2)}\right]\quad=\quad f\left[\color{royalblue}{2}\right]\quad=\quad -4\]

zepdrix (zepdrix):

\[\Large\bf\sf k'(2)\quad=\quad f'\left[\color{royalblue}{g(2)}\right]\cdot g'(2)\] \[\Large\bf\sf =\quad f'\left[\color{royalblue}{2}\right]\cdot g'(2)\quad=\quad 3\cdot 5\]